# Bayesian model-data comparison incorporating theoretical uncertainties

### v2

@@ Abstract
Accurate comparisons between theoretical models and experimental data are critical for scientific progress. However, inferred physical model parameters can vary significantly with the chosen physics model, highlighting the importance of properly accounting for theoretical uncertainties. In this article, we explicitly incorporate these uncertainties using Gaussian processes that model the domain of validity of theoretical models, integrating prior knowledge about where a theory applies and where it does not. We demonstrate the effectiveness of this approach using two systems: a simple ball drop experiment and multi-stage heavy-ion simulations. In both cases incorporating model discrepancy leads to improved parameter estimates, with systematic improvements observed as additional experimental observables are integrated.
I Introduction
@@
Comparisons between theoretical models and experimental data are at the heart of scientific inquiry. Theoretical models guide our understanding of complex systems by translating hypotheses into quantitative predictions that can be tested experimentally. Traditionally, a close fit between a model’s predictions and measured data is interpreted as a sign of success, often implying that the model parameters capture the underlying physical processes. However, this paradigm assumes that the model fully represents the complexity of actual systems – an assumption that is rarely justified in practice. All models have inherent limitations beyond their domains of validity, and using them beyond these regimes without accounting for theoretical uncertainties can lead to biased parameter estimates, reducing these parameters to mere “fitting variables” rather than meaningful physical quantities [1]. Moreover, discrepancies between certain measurements and otherwise successful models sometimes lead researchers to assign lower weights to these data, thereby diminishing their utility and limiting the potential insights they can provide.
@@
Quantifying theoretical uncertainties remains a significant challenge, particularly in systems with intricate interactions or multistage dynamics. A striking example are relativistic heavy-ion collisions where nuclei collide at near-light speeds and produce quark-gluon plasma (QGP) [2, 3]. The evolution of this deconfined matter involves several stages — initial energy deposition, hydrodynamic expansion, hadronization, and freeze-out — each introducing additional layers of assumptions and uncertainties into the model [4, 5, 6]. Without a systematic framework to disentangle true physical effects from model inadequacies, parameter extraction risks leading to overfitting rather than genuine discovery.
@@
A model discrepancy framework employing Gaussian processes (GPs) was introduced in [1], and variations have been explored in many studies [7, 8, 9, 10, 11, 12]. In these approaches, the discrepancy between experimental data and theoretical model predictions, stemming from missing physics or approximations, is modeled using a GP. However, one persistent challenge is the need to constrain the GP’s covariance kernel. For example, in [11], the authors emphasized the importance of incorporating knowledge of the theory’s validity at specific points in the input space (i.e., the domain in which observables are measured) so that both the GP and its derivative could be accurately constrained. In practice, however, specifying such accurate knowledge about the theory is often difficult. In this article, we construct the GP covariance kernel based on only qualitative prior knowledge of the theory’s domain of validity across the input space. This type of knowledge – for example, recognizing that “the theory is more reliable in this regime than in that one” – is typically easier to provide and often available. By leveraging this information, the framework prioritizes the accurate extraction of model parameters rather than simply optimizing the fit to the observables. We perform Bayesian parameter inference to simultaneously estimate both the model parameters and the GP hyperparameters, thereby quantifying uncertainties from both the experimental data and the theoretical model.
@@
We test the framework on two systems, a motivating ball-drop experiment and various multi-stage simulations of heavy-ion collisions. By comparing our results with those obtained using standard Bayesian inference, which does not account for theoretical uncertainties, we show that incorporating a constrained model discrepancy into the statistical analysis yields parameter estimates that are more robust and closer to the true values. Furthermore, the accuracy of the inferred model parameters is shown to improve systematically when increasing the number of experimental observables.
II Framework
@@
We begin by setting up the model discrepancy framework following [1]. Let the $i$-th measurement for an observable $y$ be denoted as
$y(x_{i})\sim\zeta(x_{i})+\epsilon_{i}\,,\qquad\epsilon_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})\,.$ (1)
@@
here $\delta(x)$ quantifies the discrepancy between the theoretical model prediction and the true system value at input setting $x$. Combining the above equations (1) and (2), we express the observation $y(x_{i})$ as
$y(x_{i})\sim\eta(x_{i},\bm{\theta})+\delta(x_{i})+\epsilon_{i}\,.$ (3)
@@
Following Kennedy and O’Hagan [1], we represent $\delta(x)$ as a zero-mean Gaussian process (GP) [13]:
@@
where $K(\cdot,\cdot\mid\bm{\phi})$ is the covariance kernel, which depends on a set of hyperparameters $\bm{\phi}$. A GP represents a distribution over functions, with the covariance kernel $K(\cdot,\cdot\mid\bm{\phi})$ serving as the central object that defines the properties of the functions in the distribution. In this framework, we incorporate prior knowledge about the theoretical model’s uncertainties through the choice of the covariance kernel. Adopting a Bayesian approach, we assign prior probability distributions to both the model parameters $\bm{\theta}$ and the GP hyperparameters $\bm{\phi}$, and update these to obtain their posterior distributions conditioned on the observations. In all examples discussed in this work, uniform priors (in appropriate ranges) are used for both the model parameters and the GP hyperparameters.
II.1 Choice of covariance kernel
In the examples explored in this work, we consider two different covariance kernels for the discrepancy GP, each encoding distinct prior knowledge about our confidence in the model’s validity across different values of $x$ [^2] The choice of covariance kernel for the discrepancy GP should reflect prior knowledge about theory error. In the examples considered here, we assume that only the validity of the theory as a function of the input space is known, and thus the kernels are selected accordingly. If more detailed information about the theory error is available, a more precise kernel choice can be made to incorporate this knowledge explicitly.
@@ II.1 Choice of covariance kernel
    This kernel encodes the prior belief that the model is more reliable at small $x$ than at large $x$. The term $s^{2}$ represents a baseline variance that is present across all $x$. The power term $(x_{i}x_{j})^{r}$ increases with $x$ (for $r\geq 0$ to ensure a positive definite real covariance matrix), amplifying the overall magnitude of the covariance at larger $x$. This reflects increasing uncertainty in the model at large $x$, incorporating our prior belief. The exponential decay term $\exp(-\frac{\lVert x_{i}-x_{j}\rVert^{2}}{2\ell^{2}})$ ensures that discrepancies remain correlated over short distances in $x$, but correlations decay as $x_{i}$ and $x_{j}$ become further apart. This controls how quickly the model discrepancy varies with respect to changes in $x$ and adds flexibility to the modeling. Finally, the parameter $\bar{c}^{2}$ scales the overall magnitude of the second term, determining the strength of the discrepancy relative to the baseline term $s^{2}$. This kernel has four hyperparameters: $\bm{\phi}=(s,\bar{c},r,\ell)$.
@@
Figure 1: Results of Bayesian inference for the ball drop experiment. Top row: Corner plots displaying the inferred posteriors for $g$ and $v_{0}$ (with median ±$\pm 68\%$ CI displayed on the diagonals) as additional observables (specified above the panels) are sequentially incorporated (from left to right) into the Bayesian inference. Bottom row: Model predictions based on the inferred parameters corresponding to the cases in the top row.
III Ball drop experiment
We test the model discrepancy framework first on a simple system in which a ball is dropped from a tower of height $h_{0}=60\,$m at time $t=0\,$[14]. Measurements of the ball’s height above the ground, velocity, and acceleration are recorded at discrete time points until $t=1$ s. [^3] In this example $x=t$/(1 s) serves as the unitless input parameter, $0\leq x\leq 1$. The true theory, used to generate mock experimental data, incorporates a gravitational force $\mathbf{f}_{G}=-Mg\,\hat{\bf z}$ and a drag force given by $\mathbf{f}_{D}=-0.4v^{2}\,\hat{\bf{v}}$. Here, $M=1\,$kg is the mass of the ball, $g=9.8\,{\rm m/s}^{2}$ the acceleration due to gravity, and we set the initial velocity $v_{0}$ to $v_{0}=0\,$m/s.
@@ III Ball drop experiment
Figure 2: Results for the two-parameter hydrodynamic simulation using mock data generated with fixed $\eta/s=0.1$ and $\epsilon_{\rm sw}=0.2\,$GeV/fm3. Top row: Corner plots showing the posteriors for the inferred $\eta/s$ and $\epsilon_{\rm sw}$ (with median ±$\pm 68\%$ CI displayed on the diagonals) as more observables are included in the Bayesian inference (from left to right). Bottom rows: Model predictions based on inference using all observables.
@@
Figure 3: Results for the five-parameter hydrodynamic simulation with parametrized $\eta/s$ using mock data generated with fixed $\eta/s=0.1$ and $\epsilon_{\rm sw}=0.2\,$GeV/fm3. Top row: Plots display the $\eta/s$ posterior (median and ±$\pm 95\%$ CI) as a function of temperature, with the posterior for $\epsilon_{\rm sw}$ (median ±$\pm 68\%$ CI) shown in the inset, as additional observables are sequentially incorporated into the Bayesian inference (from left to right). Bottom rows: Model predictions based on inference using all observables.
[…]
IV Heavy-ion simulation
In the following examples we simulate collisions of Au on Au nuclei at $\sqrt{s_{\textrm{NN}}}=200\,$GeV at $20-30\%$ collision centrality, using the iEBE-MUSIC simulation framework [15, 16]. Initial conditions are generated using the MC-Glauber model and coupled with relativistic viscous hydrodynamics MUSIC [17] assuming longitudinal boost-invariance, followed by hadronic transport using the UrQMD model [18, 19]. When the energy density in a cell drops below the value $\epsilon_{\rm sw}$, the fluid is “particlized”, using the iS3D hadron sampler [15] which converts the fluid’s energy and momentum into various hadron species $i$ with momenta ${\bf p}{\,=\,}({\bf p_{T}},p_{L})$ using the Cooper-Frye prescription [20]:
@@ IV Heavy-ion simulation
Here $\Sigma$ is the particlization hypersurface [21], characterized by a constant energy density $\epsilon_{\rm sw}$, $g_{i}$ is the spin degeneracy factor for particle species $i$, $f_{\mathrm{eq}}$ stands for the thermal equilibrium distribution and $\delta f$ is the out-of-equilibrium corrections.
In the following study these simulations are considered to represent the “truth”. We use them to generate mock experimental data by making measurements of the charged hadron multiplicity $E(dN_{\text{ch}}/d^{3}p)$, their elliptic and triangular flows $v_{2}\{2\}_{\rm ch}$ and $v_{3}\{2\}_{\rm ch}$, as well as the multiplicities $E(dN/d^{3}p)$ and elliptic flows $v_{2}\{2\}$ for identified pions and protons, in ten $p_{T}$ bins spanning the range $0{-}2\,$GeV. [^4] In this example the scaled transverse momentum $x=p_{T}$/(1 GeV) serves as the input parameter, $0\leq x\leq 2$. In the “truth” model we use in the particlization step the Grad 14-moment momentum distribution [22] for $f_{\mathrm{eq}}+\delta f$, and we set the switching energy density to $\epsilon_{\rm sw}=0.2\,$GeV/fm3. For various different examples, we consider $\eta/s$ either as a constant, or as a function of temperature parametrized by [23]
@@
IV.1 Truth: constant $\eta/s$; two model parameters.
[…]
@@
IV.2 Truth: constant $\eta/s$; five model parameters.
[…]
When only the momentum distribution $E(dN_{\text{ch}}/d^{3}p)$ is considered (top left), all three cases yield posteriors that are statistically compatible with the truth. However, for $\eta/s$ the posteriors accounting for MD (blue and green) overlap significantly better with the true values than those without MD (red). Once $v_{2}\{2\}_{\rm ch}$ and $v_{3}\{2\}_{\rm ch}$ are added to the calibration data (top middle), none of the posteriors uniformly covers the truth any more. When not accounting for MD (red), the posterior credible intervals for $\eta/s$ are very narrow (suggesting high confidence) but are completely inconsistent with the truth. Including MD helps a little but still infers $95\%$ credible intervals that do not cover the true value of $\eta/s$ over most of the temperature range. Apparently, the inference algorithm tries to compensate for the model’s tendency to over-predict $v_{2}$ and $v_{3}$ by increasing $\eta/s$ [^5] This inability of the MD framework to accurately identify the contribution of model parameters and GP hyperparameters, respectively, has been discussed in the model discrepancy literature [1, 7, 8, 9, 10, 11, 12]. We note that the framework presented in this work for selecting the kernel differs from earlier approaches, thereby altering the impact of this identifiability issue. When all observables (including the spectra and elliptic flows of identified pions and protons) are added to the calibration data (top right), all posteriors become narrower (i.e., more constrained). The overlap of the posteriors accounting for MD improves, whereas the posteriors that do not account for MD degrade further.
@@
IV.3 Truth: temperature-dependent $\eta/s$; five model parameters.
[…]
@@
V Summary and Outlook
Often the goal of model-data comparison is to extract physically meaningful model parameters, not solely to fit data. In this work, we introduce a framework that explicitly incorporates theoretical uncertainties by using Gaussian processes to encode prior information about a theory’s domain of validity. By integrating prior knowledge about where a theory applies and where it does not, our approach yields more robust and precise parameter estimates, as demonstrated in two different systems. In each case, systematic improvements are observed as additional experimental observables are incorporated into the calibration, leading to results that significantly outperform conventional Bayesian inference without model discrepancy (MD).
The MD framework presented in this article opens a new avenue for robust model-data comparison. With readily available qualitative information about a theory’s domain, the framework can be used to obtain reliable physical parameter estimates while correctly assessing the model’s limitations and avoiding overfitting. Importantly, it allows for the extraction of meaningful information from data that are known a priori to differ from the theoretical model — data that are typically discarded in conventional analyses. It also allows models of varying complexity and domains of validity to be directly compared against data. Finally, the framework enables validation of the inferred parameter posteriors in two ways: (i) posteriors should remain consistent as more observables are included in inference, and (ii) results obtained using a kernel with strong prior information (e.g., Kernel II) should be tighter yet consistent with those derived from more conservative modeling assumptions (e.g., Kernel I).
Acknowledgements.
We thank Daniel Phillips for useful feedback on the manuscript. This research was supported by the CSSI program Award OAC-2004601 (BAND collaboration [24]) (S.J., R.J.F., U.H, M.T.P.). C.S. is supported by the U.S. Department of Energy (DOE), Office of Science, Office of Nuclear Physics, under DOE Award No. DE-SC0021969 and DE-SC0024232. C.S. acknowledges a DOE Office of Science Early Career Award. R.J.F. was supported in part by the National Science Foundation under award PHY-2209442. R.J.F. also acknowledges support from the ExtreMe Matter Institute EMMI at the GSI Helmholtzzentrum für Schwerionenforschung GmbH, Darmstadt, Germany. Heavy-ion simulations were done using services provided by the OSG Consortium [25, 26, 27, 28], which is supported by the National Science Foundation awards #$\#2030508$ and #$\#2323298$. Bayesian inference was done on resources provided by the Ohio Supercomputer Center [29] (Project No. PAS0254).
Appendix A Corner plots for model parameters

### v3

@@ Abstract
Accurate comparisons between theoretical models and experimental data are critical for scientific progress. However, inferred physical model parameters can vary significantly with the chosen physics model, highlighting the importance of properly accounting for theoretical uncertainties. In this Letter, we present a Bayesian framework that explicitly quantifies these uncertainties by statistically modeling theory errors, guided by qualitative knowledge of a theory’s varying reliability across the input domain. We demonstrate the effectiveness of this approach using two systems: a simple ball drop experiment and multi-stage heavy-ion simulations. In both cases incorporating model discrepancy leads to improved parameter estimates, with systematic improvements observed as additional experimental observables are integrated.
journal: Physics Letters B
1 Introduction
@@
Comparisons between theoretical models and experimental data are at the heart of scientific inquiry. Theoretical models guide our understanding of complex systems by translating hypotheses into quantitative predictions that can be tested experimentally. Traditionally, a close fit between a model’s predictions and measured data is interpreted as a sign of success, often implying that the model parameters capture the underlying physical processes. However, this paradigm assumes that the model fully represents the complexity of actual systems – an assumption that is rarely justified in practice. All models have inherent limitations beyond their domains of validity, and using them beyond these regimes without accounting for theoretical uncertainties can lead to biased parameter estimates, reducing these parameters to mere “fitting variables” rather than meaningful physical quantities Kennedy2002 . Moreover, discrepancies between certain measurements and otherwise successful models sometimes lead researchers to assign lower weights to these data, thereby diminishing their utility and limiting the potential insights they can provide.
@@
Quantifying theoretical uncertainties remains a significant challenge, particularly in systems with intricate interactions or multistage dynamics. A striking example are relativistic heavy-ion collisions where nuclei collide at near-light speeds and produce quark-gluon plasma (QGP) Gyulassy:2004zy , Harris:2023tti . The evolution of this deconfined matter involves several stages — initial energy deposition, hydrodynamic expansion, hadronization, and freeze-out — each introducing additional layers of assumptions and uncertainties into the model Gale:2013da , Shen:2020mgh , Heinz:2024jwu . Without a systematic framework to disentangle true physical effects from model inadequacies, parameter extraction risks leading to overfitting rather than genuine discovery.
@@
A model discrepancy framework employing Gaussian processes (GPs) was introduced in Kennedy2002 , and variations have been explored in many studies Higdon2004 , Bayarri01052007 , Arendt:etal2012:1 , Arendt:etal2012:2 , Brynjarsdottir_2014 , GARDNER2021107381 . In these approaches, the discrepancy between experimental data and theoretical model predictions, stemming from missing physics or approximations, is modeled using a GP. However, one persistent challenge is the need to constrain the GP’s covariance kernel. For example, in Brynjarsdottir_2014 , the authors emphasized the importance of incorporating knowledge of the theory’s validity at specific points in the input space (i.e., the domain in which observables are measured) so that both the GP and its derivative could be accurately constrained. In practice, however, specifying such accurate knowledge about the theory is often difficult. In this Letter, we construct the GP covariance kernel based on only qualitative prior knowledge of the theory’s domain of validity across the input space. This type of knowledge – for example, recognizing that “the theory is more reliable in this regime than in that one” – is typically easier to provide and often available. By leveraging this information, the framework prioritizes the accurate extraction of model parameters rather than simply optimizing the fit to the observables. We perform Bayesian parameter inference to simultaneously estimate both the model parameters and the GP hyperparameters, thereby quantifying uncertainties from both the experimental data and the theoretical model.
@@
We test the framework on two systems, a motivating ball-drop experiment and various multi-stage simulations of heavy-ion collisions. By comparing our results with those obtained using Bayesian inference that does not account for theoretical uncertainties, we show that incorporating a constrained model discrepancy into the statistical analysis yields parameter estimates that are more robust and closer to the true values. Furthermore, the accuracy of the inferred model parameters is shown to improve systematically when increasing the number of experimental observables.
2 Framework
@@
We begin by setting up the model discrepancy framework following Kennedy2002 . Let the $i$-th measurement for an observable $y$ be denoted as
$y(x_{i})=\zeta(x_{i})+\epsilon_{i}\,,\qquad\epsilon_{i}\sim\mathcal{N}(0,\sigma_{i}^{2})\,.$ (1)
@@
here $\delta(x)$ quantifies the discrepancy between the experimental data and the theoretical model prediction with the true model parameters at input setting $x$. Combining the above equations (1) and (2), we express the observation $y(x_{i})$ as
$y(x_{i})=\eta(x_{i},\bm{\theta})+\delta(x_{i})+\epsilon_{i}\,.$ (3)
@@
If the functional form of $\delta(x)$ were known exactly, fitting the corrected model $\eta(x,\bm{\theta})+\delta(x)$ to data would, in principle, yield correct estimates of $\bm{\theta}$. However, $\delta(x)$ is usually not known precisely, and at best only qualitative knowledge about this error is known. We therefore treat $\delta(x)$ statistically and model it with a Gaussian process that incorporates this knowledge of the theory error.
Following Kennedy and O’Hagan Kennedy2002 , we represent $\delta(x)$ as a zero-mean Gaussian process (GP) Rasmussen2004 :
@@
where $K(\cdot,\cdot\mid\bm{\phi})$ is the covariance kernel, which depends on a set of hyperparameters $\bm{\phi}$. A GP represents a distribution over functions, with the covariance kernel $K(\cdot,\cdot\mid\bm{\phi})$ serving as the central object that defines the properties of the functions in the distribution. In this framework, we incorporate prior knowledge about the error of the theory through the choice of the covariance kernel. This prior information mitigates the identifiability problem between the theoretical model and the model discrepancy, as discussed in detail in Ref. Brynjarsdottir_2014 . Adopting a Bayesian approach, we assign prior probability distributions to both the model parameters $\bm{\theta}$ and the GP hyperparameters $\bm{\phi}$, and update these to obtain their posterior distributions conditioned on the observations. In all examples discussed in this work, uniform priors (in appropriate ranges) are used for both the model parameters and the GP hyperparameters.
2.1 Choice of covariance kernel
In the examples explored in this work, we consider two different covariance kernels for the discrepancy GP, each encoding distinct prior knowledge about our confidence in the model’s validity across different values of xx. [^2] The choice of covariance kernel for the discrepancy GP should reflect prior knowledge about theory error. In the examples considered here, we assume that the only information known about the theory error is that it increases as a function of the input space, and thus the kernels are constructed accordingly. If more detailed information about the theory error is available, a more precise kernel can be constructed to incorporate this knowledge explicitly.
@@ 2.1 Choice of covariance kernel
    This kernel encodes the prior belief that the model is more reliable at small $x$ than at large $x$. The term $s^{2}$ represents a baseline variance that is present across all $x$. The power term $(x_{i}x_{j})^{r}$ increases with $x$ (for $r\geq 0$), amplifying the magnitude of the variance at larger $x$. This reflects increasing uncertainty in the model at large $x$, incorporating our prior belief about the theory error. The exponential decay term $\exp(-\frac{\lVert x_{i}-x_{j}\rVert^{2}}{2\ell^{2}})$ ensures that discrepancies remain correlated over short distances in $x$, but correlations decay as $x_{i}$ and $x_{j}$ become further apart. Finally, the parameter $\bar{c}^{2}$ scales the overall magnitude of the second term, determining the strength of the discrepancy relative to the baseline term $s^{2}$. This kernel has four hyperparameters: $\bm{\phi}=(s,\bar{c},r,\ell)$.
@@
Figure 1: Results of Bayesian inference for the ball drop experiment. Top row: Corner plots displaying the inferred posteriors for $g$ and $v_{0}$ (with median $\pm 68\%$ CI displayed on the diagonals) as additional observables (specified above the panels) are sequentially incorporated (from left to right) into the Bayesian inference. Bottom row: Model predictions based on the inferred parameters corresponding to the cases in the top row.
@@
3 Ball drop experiment
We test the model discrepancy framework first on a simple system in which a ball is dropped from a tower of height $h_{0}=60\,$m at time t=0t=0\,Higdon:etal:2010 . Measurements of the ball’s height above the ground, velocity, and acceleration are recorded at discrete time points until $t=1$ s. [^3] In this example x=tx=t/(1 s) serves as the unitless input parameter, $0\leq x\leq 1$. The true theory, used to generate mock experimental data, incorporates a gravitational force $\mathbf{f}_{G}=-Mg\,\hat{\bf z}$ and a drag force given by $\mathbf{f}_{D}=-0.4v^{2}\,\hat{\bf{v}}$. Here, M=1M=1\,kg is the mass of the ball, $g=9.8\,{\rm m/s}^{2}$ the acceleration due to gravity, and we set the initial velocity $v_{0}$ to $v_{0}=0\,$m/s.
@@ 3 Ball drop experiment
Figure 2: Predictions based on inferred parameters when all observables are considered in calibration. The black curve shows the prediction from the true theory. The blue band shows the predictions of model + discrepancy GP with Kernel I, and green shows the same with kernel II. The red band shows the model predictions without accounting for model discrepancy (same as in the bottom-right panel of Fig. 1)
@@
We show the model + discrepancy GP predictions in Fig. 2, computed using the inferred model and GP parameters with all observables included in the calibration. The blue band shows the combined predictions with Kernel I, and the green band shows those with Kernel II. In both cases, the combined predictions agree with the observations, as expected from Eq. (3), provided the discrepancy GP is sufficiently flexible to model the theoretical error. The predictions for velocity and acceleration remain consistent with the correct model (black curve) in extrapolation regions. For height, the predictions deviate from the correct model immediately outside the range of the observations. The uncertainties of the combined predictions for all observables increase rapidly in extrapolation regions because the discrepancy GP kernels include a variance envelope proportional to $(x_{i}x_{j})^{r}$. A more theory-error informed kernel may extrapolate better. Extrapolation issues in the model discrepancy framework are discussed in detail in Ref. Brynjarsdottir_2014 .
4 Heavy-ion simulation
Figure 3: Results for the two-parameter hydrodynamic simulation using mock data generated with fixed $\eta/s=0.1$ and $\epsilon_{\rm sw}=0.2\,$GeV/fm3. Top row: Corner plots showing the posteriors for the inferred $\eta/s$ and $\epsilon_{\rm sw}$ (with median $\pm 68\%$ CI displayed on the diagonals) as more observables are included in the Bayesian inference (from left to right). Bottom rows: Model predictions based on inference using all observables.
In the following examples we simulate collisions of Au on Au nuclei at $\sqrt{s_{\textrm{NN}}}=200\,$GeV at $20-30\%$ collision centrality, using the iEBE-MUSIC simulation framework Shen:2014vra , Shen:2022oyg . Initial conditions are generated using the MC-Glauber model and coupled with relativistic viscous hydrodynamics MUSIC Paquet:2015lta assuming longitudinal boost-invariance, followed by hadronic transport using the UrQMD model Bass:1998ca , Bleicher:1999xi . When the energy density in a cell drops below the value $\epsilon_{\rm sw}$, the fluid is “particlized”, using the iS3D hadron sampler Shen:2014vra which converts the fluid’s energy and momentum into various hadron species $i$ with momenta ${\bf p}{\,=\,}({\bf p_{T}},p_{L})$ using the Cooper-Frye prescription Cooper:1974mv :
@@ 4 Heavy-ion simulation
Here $\Sigma$ is the particlization hypersurface Huovinen:2012is , characterized by a constant energy density $\epsilon_{\rm sw}$, $g_{i}$ is the spin degeneracy factor for particle species $i$, $f_{\mathrm{eq}}$ stands for the thermal equilibrium distribution and $\delta f$ is the out-of-equilibrium corrections.
In the following study these simulations are considered to represent the “truth”. We use them to generate mock experimental data by making measurements of the charged hadron multiplicity $E(dN_{\text{ch}}/d^{3}p)$, their elliptic and triangular flows $v_{2}\{2\}_{\rm ch}$ and $v_{3}\{2\}_{\rm ch}$, as well as the multiplicities $E(dN/d^{3}p)$ and elliptic flows $v_{2}\{2\}$ for identified pions and protons, in ten $p_{T}$ bins spanning the range 020{-}2\,GeV. [^4] In this example the scaled transverse momentum $x=p_{T}$/(1 GeV) serves as the input parameter, $0\leq x\leq 2$. For inference, each $p_{T}$ bin is represented by its midpoint value. In the “truth” model we use in the particlization step the Grad 14-moment momentum distribution Denicol:2012cn for $f_{\mathrm{eq}}+\delta f$, and we set the switching energy density to $\epsilon_{\rm sw}=0.2\,$GeV/fm3. For various different examples, we consider $\eta/s$ either as a constant, or as a function of temperature parametrized by JETSCAPE:2020mzn
@@
Figure 4: Results for the five-parameter hydrodynamic simulation with parametrized $\eta/s$ using mock data generated with fixed $\eta/s=0.1$ and $\epsilon_{\rm sw}=0.2\,$GeV/fm3. Top row: Plots display the $\eta/s$ posterior (median and $\pm 95\%$ CI) as a function of temperature, with the posterior for $\epsilon_{\rm sw}$ (median $\pm 68\%$ CI) shown in the inset, as additional observables are sequentially incorporated into the Bayesian inference (from left to right). Bottom rows: Model predictions based on inference using all observables.
4.1 Truth: constant $\eta/s$; two model parameters.
In this example, mock experimental data are generated using a constant $\eta/s{\,=\,}0.1$ and $\epsilon_{\rm sw}{\,=\,}0.2\,$GeV/fm3. From these data we infer the two model parameters $\eta/s$ and $\epsilon_{\rm sw}$ using a model that ignores viscous corrections during particlization.
@@
Figure 5: Results for the five-parameter hydrodynamic simulation with parametrized $\eta/s$ using mock data generated using a parametrization of $\eta/s$ with $T_{\rm kink}{\,=\,}0.18\,{\rm GeV}$, $a_{\rm low}{\,=\,}{-}1\,{\rm GeV}^{-1}$, $a_{\rm high}{\,=\,}1\,{\rm GeV}^{-1}$, and $(\eta/s)_{\rm kink}{\,=\,}0.1$, along with $\epsilon_{\rm sw}{\,=\,}0.2\,$GeV/fm3. The plot layouts and legends are identical to those in Fig. 4. The corner plot for all five model parameters, when all available observables are included, is provided in A, Fig. 7 .
4.2 Truth: constant $\eta/s$; five model parameters.
Using the same mock experimental data as before we now try to infer from them both $\epsilon_{\rm sw}$ and the temperature dependence of $\eta/s$, using a parametrization of $(\eta/s)(T)$ with four parameters ($T_{\rm kink}$, $a_{\rm low}$, $a_{\rm high}$, and $(\eta/s)_{\rm kink}$) taken from JETSCAPE:2020mzn . The true value of $\eta/s$ is 0.1, without any dependence on $T$. Note that this corresponds to a singular line in the 4-dimensional parameter space at $(\eta/s)_{\rm kink}{\,=\,}0.1$, $a_{\rm low}{\,=\,}a_{\rm high}{\,=\,}0$, and $T_{\rm kink}$ undetermined. As shown in Figure 4, which we will now discuss, this last point is problematic for Bayesian parameter inference. Similar to Fig. 3, the top row displays the inferred $\eta/s$ as a function of temperature and (in the inset) the posterior for $\epsilon_{\rm sw}$, with the same color coding as before. The corner plot for all five model parameters, for the case when all available observables are included, is provided in Fig. 6 in A. The bottom row of panels for this 5-parameter model mirrors those in Fig. 3 for the 2-parameter model.
When only the momentum distribution $E(dN_{\text{ch}}/d^{3}p)$ is considered (top left), all three cases yield posteriors that are statistically compatible with the truth. However, for $\eta/s$ the posteriors accounting for MD (blue and green) overlap significantly better with the true values than those without MD (red). Once $v_{2}\{2\}_{\rm ch}$ and $v_{3}\{2\}_{\rm ch}$ are added to the calibration data (top middle), none of the posteriors uniformly covers the truth any more. When not accounting for MD (red), the posterior credible intervals for $\eta/s$ are very narrow (suggesting high confidence) but are completely inconsistent with the truth. Including MD helps a little but still infers $95\%$ credible intervals that do not cover the true value of $\eta/s$ over most of the temperature range. Apparently, the inference algorithm tries to compensate for the model’s tendency to over-predict $v_{2}$ and $v_{3}$ by increasing $\eta/s$. [^5] This inability of the MD framework to uniquely identify the contribution of model parameters and GP hyperparameters, respectively, has been discussed in the model discrepancy literature Kennedy2002 , Higdon2004 , Bayarri01052007 , Arendt:etal2012:1 , Arendt:etal2012:2 , Brynjarsdottir_2014 , GARDNER2021107381 . We note that the framework presented in this work for selecting the kernel differs from earlier approaches, thereby altering the impact of this identifiability issue. When all observables (including the spectra and elliptic flows of identified pions and protons) are added to the calibration data (top right), all posteriors become narrower (i.e., more constrained). The overlap of the posteriors accounting for MD improves, whereas the posteriors that do not account for MD degrade further.
@@
4.3 Truth: temperature-dependent $\eta/s$; five model parameters.
In this example the mock experimental data are generated using a temperature dependent $(\eta/s)(T)$ JETSCAPE:2020mzn , with parameters set to $T_{\rm kink}{\,=\,}0.18\,{\rm GeV}$, $a_{\rm low}{\,=\,}{-}1\,{\rm GeV}^{-1}$, $a_{\rm high}{\,=\,}1\,{\rm GeV}^{-1}$, and $(\eta/s)_{\rm kink}{\,=\,}0.1$. The model remains the same as in the preceding case, with a total of five parameters.
@@
5 Summary and Outlook
Often the goal of model-data comparison is to extract physically meaningful model parameters, not solely to fit data. In this work, we introduce a framework that explicitly quantifies theoretical uncertainties by statistically modeling theory errors using Gaussian processes, guided by qualitative knowledge of a theory’s varying reliability across the input domain. This approach yields more robust and precise parameter estimates, as demonstrated in two different systems. In each case, systematic improvements are observed as additional experimental observables are incorporated into the calibration, leading to results that significantly outperform conventional Bayesian inference without model discrepancy (MD).
The MD framework presented in this Letter opens a new avenue for robust model-data comparison. With readily available qualitative information about a theory’s domain, the framework can be used to obtain reliable physical parameter estimates while correctly assessing the model’s limitations and avoiding overfitting. Importantly, it allows for the extraction of meaningful information from data that are known a priori to differ from the theoretical model — data that are typically discarded in conventional analyses. It also allows models of varying complexity and domains of validity to be directly compared against data. Finally, the framework enables validation of the inferred parameter posteriors in two ways: (i) posteriors should remain consistent as more observables are included in inference, and (ii) results obtained using a kernel with strong prior information (e.g., Kernel II) should be tighter yet consistent with those derived from more conservative modeling assumptions (e.g., Kernel I).
Data availability
The code and data to reproduce these results is publicly available in jaiswal_2025_16986775 .
Acknowledgements
We thank Daniel Phillips for useful feedback on the manuscript. S.J. thanks Jean-Paul Blaizot, Chandrodoy Chattopadhyay, Yi Chen, Sangyong Jeon, and Jean-Yves Ollitrault for insightful discussions and feedback. This research was supported by the CSSI program Award OAC-2004601 (BAND collaboration BAND_Framework ) (S.J., R.J.F., U.H, M.T.P.). C.S. is supported by the U.S. Department of Energy (DOE), Office of Science, Office of Nuclear Physics, under DOE Award No. DE-SC0021969 and DE-SC0024232. C.S. acknowledges a DOE Office of Science Early Career Award. R.J.F. was supported in part by the National Science Foundation under award PHY-2209442. R.J.F. also acknowledges support from the ExtreMe Matter Institute EMMI at the GSI Helmholtzzentrum für Schwerionenforschung GmbH, Darmstadt, Germany. Heavy-ion simulations were done using services provided by the OSG Consortium osg07 , osg09 , osg77 , osg78 , which is supported by the National Science Foundation awards #2030508\#2030508 and #2323298\#2323298. Bayesian inference was done on resources provided by the Ohio Supercomputer Center OhioSupercomputerCenter1987 (Project No. PAS0254).
Appendix A Corner plots of model parameters for heavy-ion simulation